OpenStudy (anonymous):

Please HELP I will give you a medal :) Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

I know that the next step is to add 2 to each side right? But which one shows that?

OpenStudy (anonymous):

I think It's -x-2=4 and x-2=-4 is this 6th grade math?

OpenStudy (anonymous):

Nope 9th grade @Dalycia

OpenStudy (anonymous):

Oh Ok

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Hmm You should go with -x-2=4 But Idk sorry

OpenStudy (anonymous):

its okay here is your medal!

OpenStudy (anonymous):

ty